OpenStudy (pokemon23):

If 14 added to a certain number and the sum is multiplied by 2 the result is equal to 8 times the number decreased by 14.Find the number. If 14 added to a certain number and the sum is multiplied by 2 the result is equal to 8 times the number decreased by 14.Find the number. @Mathematics

OpenStudy (pokemon23):

Bath guy dude :) alright high five

OpenStudy (anonymous):

slap!

OpenStudy (pokemon23):

hell yeah

OpenStudy (anonymous):

(x+14)*2=(x-14)*8

OpenStudy (anonymous):

solve it from there

OpenStudy (anonymous):

(x+14)*2=(x-14)*8 2x+28=8x-112 -6x=74 x=74/-6

OpenStudy (anonymous):

srry that took a while

OpenStudy (pokemon23):

nah it alright your awesome

OpenStudy (anonymous):

thx

OpenStudy (pokemon23):

lol we both have 12 fans